NONCONSTANT GROWTH

Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $1.25 coming 3 years from today. The dividend should grow rapidly-at a rate of 35% per year-during Years 4 and 5; but after Year 5, growth should be a constant 9% per year. If the required return on Computech is 15%, what is the value of the stock today? Round your answer to the nearest cent. Do not round your intermediate calculations.

$
Answer #1

D3=1.25

D4=(1.25*1.35)=1.6875

D5=(1.6875*1.35)=2.278125

Value after year 5=(D5*Growth rate)/(Required rate-Growth rate)

=(2.278125*1.09)/(0.15-0.09)

=41.3859375

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.25/1.15^3+1.6875/1.15^4+2.278125/1.15^5+41.3859375/1.15^5

=$23.50(Approx).